# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

### v1

@@ 4.7 Difficulty Scaling
Problem difficulty in EngChain is not a subjective label but a calibrated measure determined by systematically varying three primary factors:
  • Conceptual Complexity refers to the number and integration of engineering principles required. An Easy problem typically tests a single, isolated concept like Ohm’s Law. An Intermediate problem requires the combination of two or three concepts, such as applying the principle of superposition to calculate a net electric field from multiple point charges. An Advanced problem demands the integration of multiple complex ideas, such as using numerical integration on Levenspiel plot data to determine a reactor volume.
  • Mathematical Sophistication describes the nature of the required calculations. Easy problems involve direct algebraic substitution. Intermediate problems may require basic calculus or vector addition. Advanced problems escalate to solving differential equations or applying numerical methods, such as the trapezoidal rule for integration.
  • Procedural Depth measures the length and interdependence of the reasoning chain. An Easy problem might be solvable in one or two steps. In contrast, an Advanced problem involves a long sequence of calculations where the output of each step is a critical input for the next, testing a model’s ability to maintain a coherent and accurate computational trace.
This multi-axis framework for difficulty allows EngChain to move beyond a simple easy/hard dichotomy and provides a more fine-grained analysis of an LLM’s reasoning capabilities.
@@ 5 Evaluation Framework
For a verifiable assessment of the entire reasoning process, we compute several key metrics by comparing a model’s generated solution against the ground-truth reasoning trace. The evaluation is multifaceted. It jointly evaluates final answer accuracy, the fluency and semantic quality of the generated solution, and the correctness of the multi-step reasoning chain.
@@ 5.1 Final Answer Accuracy
It is determined by comparing the model’s predicted numerical answer ($A_{pred}$) to the ground-truth answer ($A_{gt}$). A response is considered correct only if its relative error is less than a strict tolerance ($\epsilon_{final}$) of 1%, a standard appropriate for engineering precision. The calculation is formalized as follows:
$\text{Relative Error}=\left|\frac{A_{gt}-A_{pred}}{A_{gt}}\right|<\epsilon_{final}$
@@ 5.2 Textual Quality
We assess the textual quality of a generated solution, evaluating its fluency and semantic fidelity against the ground-truth text using a combination of semantic and lexical similarity scores using the most common reference based evaluation metrics (Xie et al., 2023b):
  • BERTScore (Zhang et al., 2020) measures deep semantic similarity by aligning contextual embeddings of tokens from the generated solution and the ground-truth solution. It computes precision, recall, and an F1-score to provide a robust measure of semantic equivalence, capturing nuances that lexical overlap metrics may miss.
  • ROUGE-2 (Lin, 2004) quantifies the lexical overlap by calculating the F1-score based on the number of matching adjacent word pairs (bigrams) between the generated and ground-truth texts.
Together, these scores provide a complementary assessment that reflects the model’s ability to produce solutions using appropriate and correctly structured domain-specific terminology.
5.3 Multi-Step Reasoning Verification
This is the core of our evaluation framework which assesses the entire logical chain of a model’s solution. This process moves beyond final answer accuracy by comparing each predicted reasoning step ($P_{j}$) against each ground-truth step ($GT_{i}$). A step is considered valid only if it satisfies two criteria: semantic similarity and numerical correctness.
To formalize this, we compute a combined similarity matrix, $S_{comb}$, by taking the element-wise product (Hadamard product $\circ$) of a semantic similarity matrix ($S_{sem}$) and a numerical correctness matrix ($S_{num}$).
$S_{comb}=S_{sem}\circ S_{num}$
  • Semantic Similarity ($S_{sem}$) is computed using a powerful Cross-Encoder model (cross-encoder/stsb-roberta-large). This model processes each ground-truth and predicted step as a single contextual pair, yielding a high-fidelity similarity score that captures nuanced semantic relationships more effectively than bi-encoder models like the initially used SentenceTransformer.
  • Numerical Correctness ($S_{num}$) is a binary matrix where an entry is 1 if the intermediate numerical answer in a predicted step matches the corresponding ground-truth step within a 2% relative error tolerance, and 0 otherwise.
From this combined matrix, we determine if a step is “matched” by checking if its maximum similarity score exceeds a predefined threshold, $\tau$. We then calculate recall and precision to evaluate the reasoning chain. Recall measures the fraction of ground-truth steps the model successfully reproduced:
$\text{Recall}=\frac{1}{|GT|}\sum_{i=1}^{|GT|}\mathbb{I}\left(\max_{j}S_{comb}[i,j]>\tau\right)$
Precision measures the fraction of the model’s generated steps that were valid and relevant:
$\text{Precision}=\frac{1}{|P|}\sum_{j=1}^{|P|}\mathbb{I}\left(\max_{i}S_{comb}[i,j]>\tau\right)$
Finally, our primary metric for reasoning quality, the Step F1-Score, is calculated as the harmonic mean of the step-level recall and precision defined above, providing a single, robust measure of procedural correctness.
$\text{Step F1-Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
5.4 Automated Error Analysis with LLM-as-a-Judge
While the Step F1-Score provides a sound quantitative measure of a model’s procedural correctness, it has a significant limitation: it can identify that a reasoning chain is flawed, but it cannot explain why. For instance, as will be discussed later, our preliminary results revealed models that achieved high final-answer accuracy despite very low reasoning scores, a classic case of being “right for the wrong reasons.” This made us realize that to diagnose these failures and determine their cause, a deeper, more qualitative analysis was required. Hence, to move beyond quantitative scores and introduce a layer of explainability, we propose a novel Automated Error Analysis system that employs a large language model as an expert evaluator, or an LLM-as-a-Judge.
This system is triggered for any ground-truth step ($GT_{i}$) that fails to find a satisfactory match among the predicted steps. A step is flagged as a “failure” if its maximum combined similarity score falls below an empirically determined threshold, $\tau=0.7$. This condition is formalized as:
$\max_{j}S_{comb}[i,j]<\tau$
Figure 5: Automated error analysis with the LLM-as-a-Judge workflow.
The overall workflow of this automated analysis process is illustrated in Figure 5. At its core, the process is rooted in carefully designed prompt engineering, where the judge model is given a system prompt defining its persona as an expert engineering professor and a set of critical rules for logical consistency. The input to the judge is a structured prompt containing the full problem context, the ground-truth step, and the model’s incorrect step. Providing this complete context is crucial for the judge to accurately classify the failure into one of the five distinct categories, which are defined in Appendix I. To ensure a reliably parsable output, the judge is instructed to return only a structured JSON object containing its classification and a detailed and scalable diagnosis of the model’s specific failure modes. The full system prompt used for the automated error analysis is provided in Appendix J.
@@ 6.1 Evaluation Model Suite
To provide a comprehensive evaluation of engineering reasoning capabilities, we evaluate models from a diverse suite. The results presented in Section 7 focus on 11 distinct models from the first category. While a full evaluation of the other categories is planned for future work, we group the complete suite of models into the following four-way taxonomy:
  1. 1.
    Frontier proprietary models, which represent the current upper bound of performance. This includes models such as GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025).
  2. 2.
    Science/Engineering-enhanced models, such as Galactica (Taylor et al., 2022), Minerva (Lewkowycz et al., 2022), ChemLLM (Zhang et al., 2024), SciLit01 (Li et al., 2025a), and SciAgent (Gao et al., 2024), are systems pre-trained or fine-tuned on vast scientific and technical corpora to improve domain-specific knowledge and complex procedural reasoning.
  3. 3.
    Math-enhanced models, such as WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024), are systems specifically trained on mathematical and symbolic corpora to enhance their quantitative reasoning capabilities.
  4. 4.
    General-purpose open models, such as LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025), provide strong, domain-agnostic foundations that are often used to create specialized variants.
This four-way taxonomy provides a structure for comparing different model architectures and training paradigms, highlighting how model scale, domain specialization, and mathematical supervision influence symbolic engineering reasoning. The detailed configurations and model sources can be found in Appendix K.
@@ 6.2 Experimental Setup
We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model. We evaluate all models under a unified decoding configuration: a low temperature of 0.2 and a maximum token limit of 4,096. We use a low temperature (T=0.2) to ensure deterministic, high-confidence outputs and a high token limit (4,096) to allow ample space for detailed, multi-step solutions without premature cut-offs. We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
You are an expert engineer. Solve the following
problem by providing a detailed, structured
solution. Use the exact headings and
formatting provided below.
# Given
List all known variables and their
values with units.
# Find
State the variable(s) to be calculated.
# Formulae
Write down all necessary governing equations
before substituting any values.
# Solution
Provide a step-by-step calculation. Each step
must start on a new line and be formatted exactly
as ’**Step X:**’, where X is the step number.
Show the substitution of values into
the formulae clearly.
# Final Answer
State the final numerical result with its units
in the format: **Answer:** [value] [units]
To process the model outputs, we first segment the text into reasoning steps. We use a regular expression to identify line-start markers that match either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.). A subsequent regex cleans these prefixes. We then extract numerical values from each step and the final answer using a three-priority heuristic: first, matching the **Answer:** tag; second, falling back to the last number following an equals sign (e.g., V = 100.0); and third, taking the last standalone number. If no overall **Answer:** tag is found, the final numerical value from the last step is used for verification.
7 Results
7.1 Overall Model Performance
The primary finding from our evaluation is a stark and universal divergence between the models’ ability to produce a correct final answer and the verifiable correctness of their reasoning process. As shown in  Table 1, while the top models achieve a final answer accuracy of approximately 63.1% (gpt-4.1-mini and gpt-5-mini), their procedural reasoning capability is critically low. The Reasoning (F1 Score), our primary metric for procedural correctness, is alarmingly poor across the board, with the top-performing model, gemini-2.5-flash, only achieving 19.32%. This demonstrates that in the vast majority of cases, all models fail to follow a sound, verifiable problem-solving methodology, confirming the “right for the wrong reasons” phenomenon at scale.
This significant reasoning gap is further masked by high textual fluency. The BERTScore, for instance, remained consistently high for all models, averaging approximately 0.87. This indicates that models excel at generating semantically appropriate, engineering-domain language, even when the underlying procedural logic is flawed or nonsensical.
Model Final Answer Accuracy $\uparrow$ Reasoning (F1 Score) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
Frontier Proprietary Models
Claude 3.7 56.8149.55 16.9121.76 88.031.16 29.287.65 37.097.63
Claude 4 57.7849.41 14.1618.28 87.931.33 29.558.16 37.257.91
Claude 4.5 52.2249.97 14.1919.72 87.391.09 28.108.15 35.347.58
Deepseek v3.2 61.2648.73 13.5319.54 86.541.23 25.698.96 34.579.12
Deepseek R1 59.9349.02 19.0324.82 86.481.34 25.588.34 34.498.30
Gemini 2.5 Flash 59.4149.13 19.3223.56 87.551.63 28.829.17 36.409.36
Gemini 2.5 Pro 59.9349.02 17.9721.01 88.241.41 31.828.29 37.918.49
GPT 4.1 57.7849.41 12.1315.85 86.651.24 26.749.13 34.459.12
GPT 4.1 Mini 63.1148.27 14.8419.69 86.521.14 25.198.17 33.828.34
GPT 5 60.4448.92 15.4721.79 88.191.21 28.988.30 36.558.06
GPT 5 Mini 63.1148.27 14.3020.32 87.951.10 26.637.62 34.747.52
Table 1: Overall performance of frontier models on the EngChain benchmark (N=1350) under a Zero-Shot setting. Scores represent the mean, with standard deviation shown in superscript. The highest mean score in each column is in bold.
7.2 Performance by Engineering Branch
A fine-grained analysis of the results reveals a significant performance disparity across the three engineering branches, as illustrated in Figure 6. We observe that model reasoning capabilities are not uniform and are highly sensitive to the specific scientific principles and mathematical structures of a given engineering domain.
Chemical Engineering consistently proved to be the most challenging domain for all models, yielding the lowest scores in both final answer accuracy and procedural reasoning. The reasoning collapse in this domain is particularly severe, with several models falling into single-digit Reasoning (F1 Score)s, including Deepseek v3.2 (6.90%) and GPT 5 Mini (7.58%).
Conversely, Mechanical Engineering was clearly the most solvable domain, where models achieved their highest scores. This branch produced the peak Final Answer Accuracy of the entire benchmark (Gemini 2.5 Pro at 70.22%) and the highest Reasoning (F1 Score) (Deepseek R1 at 30.11%).
Electrical Engineering demonstrated intermediate performance, with scores that consistently fell between the other two branches. Its peak Reasoning (F1 Score) of 22.49% (Gemini 2.5 Flash) and peak Final Answer Accuracy of 67.11% (GPT 4.1 Mini) were modest, clearly lagging behind the top scores in Mechanical Engineering but remaining well above the challenging baseline set by Chemical Engineering.
Figure 6: Branch-level performance comparison of the top four models ranked by mean Reasoning (F1 Score) across the three engineering branches. Chemical Engineering is consistently the most challenging domain for procedural reasoning.
7.3 Performance by Engineering Domain
Further analysis reveals that model performance is highly inconsistent across the nine distinct engineering domains, as visualized in the radar plot in Figure 7. This “spiky” performance profile, where a model can excel in one domain while failing in another, suggests that current models possess patches of specialized knowledge rather than generalized, first-principles reasoning.
The data shows several domains are exceptionally challenging for all models. Domains such as Reaction Kinetics, Mechanics of Materials, Transport Phenomena, and Electromagnetics & Waves produced critically low reasoning scores. In these areas, it was common for top-tier models to score in the single digits, with Deepseek v3.2 scoring 6.90% in Reaction Kinetics and GPT 5 Mini scoring 7.58% in Mechanics of Materials.
Conversely, the domain of Vibrations & Acoustics emerged as a clear outlier, representing the easiest domain for nearly all models. The reasoning scores here were dramatically higher than in any other domain, with Deepseek R1 achieving 52.26% and Gemini 2.5 Flash achieving 40.70%. The domain of Signals & Systems was another area of relative strength, where most models achieved some of their highest reasoning scores.
This domain-level variance underscores the need for broad, multi-domain benchmarks. We hypothesize this performance gap is linked to the underlying mathematical complexity. The high-scoring domains in our results often feature problems solvable with direct algebraic substitution or linear, discrete procedures. The most challenging domains, in contrast, frequently require models to set up and solve differential equations or manage systems of complex, interdependent variables, which appear to be critical failure points for current procedural reasoning.

### v2

@@ 3.4 Difficulty Scaling
Problem difficulty in EngTrace is calibrated through a multi-axis framework defined by three dimensions: Conceptual Complexity (isolated principles vs. multi-domain synthesis), Mathematical Sophistication (direct algebraic substitution vs. differential equations), and Procedural Depth (length of interdependent reasoning chains). A detailed quantitative breakdown of the dataset’s difficulty distribution is provided in Appendix K.
@@ 4 Evaluation Framework
Inspired by recent advances in process supervision (Lightman et al., 2023) and cascaded model verification (Chen et al., 2023), we introduce a two-stage evaluation framework that validates intermediate reasoning steps in addition to final-answer accuracy through a tiered verification protocol. To ensure alignment with professional standards, we validated this protocol through a blind human-expert study, confirming that our Tribunal-based approach significantly outperforms standard metrics in capturing engineering reasoning quality (see Appendix M for full methodology and ablation results).
4.1 Preliminaries
We define the evaluation of a reasoning chain as a comparison between the ground truth solution $G$ and the model-generated predicted solution $P$. Both are represented as ordered sequences of $M$ and $N$ discrete reasoning steps, respectively:
$G=(g_{1},g_{2},\dots,g_{M}),\quad P=(p_{1},p_{2},\dots,p_{N})$ (3)
Each individual reasoning step $s_{i}\in\{G,P\}$ consists of a textual derivation $t_{i}$ and, where applicable, an associated numerical or symbolic intermediate result $a_{i}=\text{Res}(s_{i})$.
To assess the validity of the derivation, we seek a mapping that aligns steps in $P$ to their logical counterparts in $G$. Unlike standard string matching, this alignment must account for structural variations such as step merging, reordering, or alternative valid methodologies while ensuring that the cumulative logic remains faithful to the underlying engineering principles.
@@ 4.2 Final Answer Accuracy
To evaluate the correctness of the final predicted outcome, let $g_{M}$ and $p_{N}$ denote the last steps of the ground truth and predicted solutions, from which we extract the terminal numerical results $A_{G}=\text{Res}(g_{M})$ and $A_{P}=\text{Res}(p_{N})$. A response is deemed correct if the relative error between these values satisfies the following condition:
$\mathbb{I}_{final}=\mathbb{I}\left(\frac{|A_{G}-A_{P}|}{|A_{G}|}<\epsilon_{final}\right)$ (4)
where $\epsilon_{final}=0.02$ (2%), a strict threshold required to prevent catastrophic error propagation in high-precision engineering derivations.
@@ 4.3 Textual Quality
We evaluate the fluency and semantic fidelity of generated solutions using standard reference-based metrics (Xie et al., 2023b). We employ BERTScore (Zhang et al., 2020) to measure deep semantic similarity via contextual embedding alignment, and ROUGE-2 (Lin, 2004) to quantify lexical overlap. Together, these metrics assess the model’s ability to employ appropriate domain-specific terminology.
4.4 Tiered Reasoning Verification
To evaluate the full reasoning chain, we characterize the validity of the predicted sequence $P$ against the ground truth $G$ through a tiered verification protocol. This process consists of two phases: determining step-level validity through hierarchical assessment and aggregating these results into a global reasoning metric.
Step-Level Validity Assessment.
We compute a validity matrix $V\in\mathbb{R}^{M\times N}$, where each element $V_{i,j}$ represents the validity of predicted step $p_{j}$ relative to ground truth step $g_{i}$. This validity is determined by two successive tiers of verification.
Tier 1: Automated Symbolic Verification.
This tier fills the matrix $V$ by enforcing syntactic and numerical precision through a binary step-validity score:
$S_{\mathcal{T}_{1}}(g_{i},p_{j})=\delta_{num}\land\delta_{sem}$ (5)
where $S_{\mathcal{T}_{1}}=1$ only if the numerical consistency (relative error tolerance $\epsilon=0.02$) and semantic similarity (Cross-Encoder threshold $\tau=0.7$) are both satisfied. To optimize computational efficiency, we calculate the alignment ratio, which represents the proportion of steps in $P$ successfully verified by Tier 1. If this ratio exceeds a confidence threshold ($\rho=0.8$) and the final answer is correct, the reasoning chain is classified as a standard derivation and accepted without further review. Chains falling below this threshold are escalated to Tier 2 for heuristic assessment.
Tier 2: Multi-Model Heuristic Verification.
Escalated chains are evaluated by a heterogeneous AI Tribunal to identify valid alternative derivations (prompt provided in Appendix L). Each judge $J_{k}$ classifies the observed discrepancy between $p_{j}$ and $g_{i}$ into a category $C_{k}$, which is then mapped to a scalar score:
$\mu(C)=\begin{cases}1.0&\text{if }C=\text{Alternate Correct}\\ 0.5&\text{if }C=\text{Calculation Error}\\ 0.0&\text{if }C\in\{\text{Conceptual Error, Other}\}\end{cases}$ (6)
The majority consensus of these scalar scores defines $S_{\mathcal{T}_{2}}$, which is assigned to $V_{i,j}$ to finalize the validity matrix for non-standard derivations. In the absence of a clear majority, the assessment defaults to a conservative tie-breaker where the lowest score is assigned.
Global Reasoning Metric $\mathbf{F1_{rec}}$.
To quantify the overall quality of the reasoning chain, we calculate the Recovered Reasoning F1 Score ($F1_{rec}$). We first identify the optimal alignment that maximizes the aggregate verification score:
$\mathcal{R}_{total}=\max_{\sigma\in\mathcal{A}}\sum_{(i,j)\in\sigma}V_{i,j}$ (7)
where $\mathcal{A}$ represents the set of valid one-to-one mappings between steps. This optimization is solved using the Hungarian matching algorithm on the finalized validity matrix $V$. We define recovered precision and recall as $P_{rec}=\mathcal{R}_{total}/N$ and $R_{rec}=\mathcal{R}_{total}/M$, respectively, where $N$ and $M$ are the number of predicted and ground truth steps. Finally, $F1_{rec}$ is computed as the harmonic mean of $P_{rec}$ and $R_{rec}$.
@@ 5.1 Evaluation Model Suite
We evaluate 24 models categorized into a three-way taxonomy to isolate the impacts of scale, access, and specialization: (1) Frontier proprietary models represent the current state-of-the-art in reasoning and instruction following. Our evaluation includes 12 models in this category: GPT-5 {Base, Mini} and GPT-4.1 {Base, Mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025c, d, a), Gemini {3 Pro, 2.5 Pro, 2.5 Flash} (Google, 2025; Comanici et al., 2025), and DeepSeek {V3, R1} (Liu et al., 2024; Guo et al., 2025a). (2) General-purpose open models provide strong, domain-agnostic foundations and allow for transparent analysis of performance relative to parameter count. We evaluate 8 models: Meta-Llama-3.1 {8B, 70B} (Grattafiori et al., 2024), Qwen2.5 {7B, 14B, 72B} and Qwen3-8B (Qwen, 2024, 2025), and Gemma {2-9B, 3-27B} (Gemma Team et al., 2024; Gemma Team and Google DeepMind, 2025). (3) Math-enhanced models are systems specifically trained on mathematical and symbolic corpora to enhance quantitative reasoning. We evaluate 4 specialized models in the 7B parameter class: Qwen2.5-Math-7B (Yang et al., 2024), Mathstral-7B (Mistral, 2024), WizardMath-7B (Luo et al., 2023), and MetaMath-7B (Yu et al., 2023). Detailed configurations and model sources can be found in Appendix N.
@@ 5.2 Experimental Setup
We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model. We evaluate all models under a unified decoding configuration: a low temperature of 0.2 to ensure deterministic outputs and a maximum token limit of 4,096 to allow for detailed reasoning. All models are evaluated in a zero-shot setting using a standardized prompt and output parsing pipeline; full implementation specifications are provided in Appendix O.
5.3 Results
5.3.1 Overall Model Performance
Table 1 summarizes performance across five key metrics: Final Answer Accuracy (FAC), Reasoning F1, and standard semantic measures (BERTScore, ROUGE-2/L). The results reveal a distinct hierarchy: frontier proprietary models define the state-of-the-art, with Gemini 3 Pro achieving peak numerical precision (FAC 64.00%) and DeepSeek R1 leading in logical process alignment (Reasoning F1 44.41%). In contrast, open-weights models lag significantly in both metrics despite competitive semantic scores, even among large-scale systems like Llama 3.1 70B. Notably, specialized math-enhanced models in the 7B parameter class generally underperform general-purpose counterparts of equivalent size, suggesting that abstract mathematical pre-training does not strictly transfer to the physically constrained reasoning required for engineering.
Model Final Answer $\uparrow$ Reasoning (F1) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
Frontier Proprietary LLMs
Claude 3.7 Sonnet 57.63 39.36 88.09 29.42 37.36
Claude 4 Sonnet 56.81 36.80 87.92 29.52 37.22
Claude 4.5 Sonnet 51.78 33.61 87.37 28.06 35.39
DeepSeek V3 62.37 37.06 86.54 25.69 34.57
DeepSeek R1 61.04 44.41 86.48 25.58 34.49
Gemini 2.5 Flash 59.78 42.22 87.64 29.14 36.70
Gemini 2.5 Pro 60.44 39.38 88.24 31.88 38.07
Gemini 3 Pro 64.00 37.05 86.71 27.32 34.99
GPT-4.1 57.11 31.22 86.63 26.78 34.46
GPT-4.1 Mini 61.56 36.87 86.50 25.06 33.87
GPT-5 61.04 41.81 88.19 28.98 36.55
GPT-5 Mini 63.70 41.01 87.95 26.63 34.74
General Purpose Open LLMs
Gemma 2 9B 26.59 21.37 88.19 28.02 35.77
Gemma 3 27B 44.30 27.59 87.04 23.93 32.66
Llama 3.1 8B 23.70 18.61 87.65 25.59 32.50
Llama 3.1 70B 40.81 30.04 88.34 28.87 35.82
Qwen 2.5 7B 35.63 23.67 86.30 24.21 32.89
Qwen 2.5 14B 42.07 27.67 86.44 24.94 33.38
Qwen 2.5 72B 48.07 28.02 86.46 25.58 33.96
Qwen 3 8B 23.70 16.68 86.91 8.93 11.09
Math Enhanced LLMs
Mathstral 7B 19.70 18.59 87.79 28.02 36.42
MetaMath 7B 6.07 2.37 82.97 10.57 16.69
Qwen 2.5 Math 7B 23.04 18.18 85.61 21.71 29.98
WizardMath 7B 8.22 6.38 83.66 14.14 20.38
Table 1: Overall Performance Comparison of 24 models on the EngTrace benchmark. Scores are reported as percentages. The Final Answer column denotes strict numerical accuracy, while Reasoning (F1) measures the structural alignment of the derivation steps. The highest score in each column is highlighted in bold.
5.3.2 Performance by Engineering Branch
Analysis by branch reveals a trade-off between reasoning fidelity and numerical precision. Chemical Engineering proved the most rigorous, yielding the lowest average scores; DeepSeek R1 achieved a decisive lead here (FAC 64.89%, F1 40.91%), confirming that complex stoichiometry demands sustained chain-of-thought capabilities. In Electrical Engineering, Gemini 3 Pro secured the top FAC (68.67%), while DeepSeek R1 retained the highest Reasoning F1 (41.22%), indicating structurally superior derivations even when numeric outputs diverge. Mechanical Engineering was the most accessible, with Gemini 2.5 Pro achieving the benchmark’s highest overall FAC (71.56%). Notably, Llama 3.1 70B showed strong semantic performance here (BERTScore 88.37%), suggesting mechanical terminology is particularly well-represented in general pre-training corpora.
5.3.3 Performance by Engineering Domain
A granular analysis of specific domains (Figure 4) reveals critical variances in capability. In Chemical Engineering, DeepSeek R1 dominates Thermodynamics and Reaction Kinetics (FAC 78.33% and 58.00%), suggesting its reinforcement learning optimization effectively navigates coupled equations; conversely, Transport Phenomena remains a bottleneck (max 53.33% FAC). Electrical Engineering results are heterogeneous: Gemini 3 Pro excels in Digital Communications (77.33% FAC), likely due to its logic-based nature, but physics-heavy domains like Electromagnetics remain challenging, with no model exceeding 56.00% FAC. In Mechanical Engineering, Mechanics of Materials appears largely solved (Gemini 3 Pro: 92.67% FAC), while Fluid Mechanics caps at 56.00% FAC. This indicates that while models have mastered static relationships, they continue to struggle with the non-linear differential equations governing fluid dynamics.